OpenStudy (anonymous):

Find the equation of the line parallel to the line y = 2x – 5. y = -2x y = 2x y = 1/2x – 5 y = -1/2x – 5

OpenStudy (saifoo.khan):

Since it says parallel, you have to find something with the "same slope".

OpenStudy (anonymous):

y = 2x

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (anonymous):

Find the equation of the line parallel to the line y = 1/4x – 2. y = -4x – 2 y = -1/4 x y = -4x y = 1/4x + 2

OpenStudy (anonymous):

for this would it be y = -4x-2?

OpenStudy (saifoo.khan):

.."parallel"..

OpenStudy (anonymous):

ohhh the last one i sorta skimmed over it :P

OpenStudy (anonymous):

is that correct tho?